A piece of wire 20 m long is cut into two pieces. One piece is bent into a square and the

other is bent into an equilateral triangle. How should the wire be cut so that the total area

enclosed is:

(a) maximum? (b) minimum?


1
Expert's answer
2021-07-26T11:19:08-0400

Let x metres be the side of the square, so 4x metres are the total amount of wire used for the square (with "0\n\u2264\n4\nx\n\u2264\n20\n\u21d2\n0\n\u2264\nx\n\u2264\n5\n )" For the equilateral triangle 20−4x metres of wire remains, and the side will be "\frac{20
−
4
x}{
3}"

Given the side l of an equilateral triangle, the height is, using Pitagora's theorem, "h
=
\frac{l}{
2}
\sqrt
3" so the area of the triangle is:

"A
=
l
⋅
\frac{l}{
2}
\sqrt
3
⋅
\frac1
2
=
l^
2
\frac{ 
{\sqrt3}}{4}"


"(
A
=
\frac{b
⋅
h}
2
)"


So the total area is:

"A=x^2+(\frac{20-4x}{3})^2 \frac{\sqrt{3}}{4}\\
A'= 2x+ \frac{\sqrt{3}}{4}*\frac{1}{9}*20(20-4x)*(-4)\\
A'=2x-2\frac{\sqrt{3}}{9}(20-4x)\\
A' > 0 \space if\\
2x-2\frac{\sqrt{3}}{9}(20-4x) > 0 \implies 18x-40\sqrt{3}+8\sqrt{3}>0 \\
x> \frac{20*3(3\sqrt{3}-4)}{81-48}\\
x> \frac{20}{11}(3\sqrt{3}-4)= \bar{x}\\"


So our function Area is decreasing in "[
0
,
\bar{x}
)" and growing in "(
\bar{x}, 5]" and so the point "(
\bar{
x}
,
A
(
\bar{x}
)
)" is a minimum of the function. The conclusion is that maximum (absolute maximum) is in one of the two sides of the interval of the definition of x, 0 or 5.

If  "x\n\n=\n\n0" the value of area will be: "A

=

100

\frac{√

3}{

9}

≅

19

.

2"

If "x\n\n=\n\n5"  the value of the area will be: "A\n\n=\n\n25" , that is greater than the other value.



Therefore

a. Minimum = 19.2

b. Maximum = 25